OpenStudy (anonymous):

Autumn is depositing $575 at the beginning of every six months into an account with an interest rate of 3.67%, compounded semiannually. She wants to renovate her kitchen in seven years. How much will Autumn have in the account at the time of the renovation?

OpenStudy (anonymous):

@mathmate @just_one_last_goodbye @Legends @leonardo0430 @DarkMoonZ

OpenStudy (mathmate):

Are you familiar with the compound interest formula?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

Can you post it?

OpenStudy (mathmate):

We will work together to get the answer.

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

A=P(1+r/n)^nt

OpenStudy (mathmate):

Excellent! Do you know what each symbol represents?

OpenStudy (anonymous):

P = 575 r = 3.67% or 0.0367 n = 2 t = 7

OpenStudy (mathmate):

Exactly!

OpenStudy (anonymous):

575(1+0.0367/2)^14 = $741.70 This, however, is not an answer choice.

OpenStudy (mathmate):

Oh, sorry! I have asked you for the compound interest formula, but since she puts in money every 6 months, the total will be much bigger. We need to use the amortization formula. Would you have that handy?

OpenStudy (anonymous):

No... :(

OpenStudy (mathmate):

Let me figure it out and post it. Give me a couple of minutes.

OpenStudy (mathmate):

ok, here it is. Let R=1+r/n=1+0.0367/2 n=14 as before

OpenStudy (mathmate):

Let P=payment for every 6 months, and A=amount at the end of 7 years. Since she puts money in the bank at the beginning of each 6-month period, so the amount she gets a the total amount is: \(A = PR + PR^2 + PR^3.... + PR^n\) ok so far?

OpenStudy (mathmate):

@MafHater ?

OpenStudy (anonymous):

Ok, so...

OpenStudy (anonymous):

P = 575

OpenStudy (anonymous):

What is R?

OpenStudy (anonymous):

interest?

OpenStudy (mathmate):

Remember we know everything on the right hand side, so in principle, we can calculate the answer.

OpenStudy (mathmate):

R=1+r/n = 1+0.0367/2=1.01835 for each period of six months.

OpenStudy (mathmate):

PR means the amount she gets from the LAST payment PR^14 means the amount she gets from the FIRST payment.

OpenStudy (mathmate):

Because the first payment has stayed in the account for 14 periods.

OpenStudy (anonymous):

Wow, I hate this. LOL

OpenStudy (mathmate):

In math, try to understand, and not memorize. Once you understand, everything clicks.

OpenStudy (anonymous):

Okay, would you mind giving me the whole equation? I will then solve it.

OpenStudy (mathmate):

BTW, the difference between evaluate and solve: If and when I give you the whole formula, you will evaluate it, because you know everything on the right hand side. When you have an equation, you solve for the unknown that is somewhere in the equation, not necessary only on the left-hand side.

OpenStudy (mathmate):

Going back to: \(A = PR + PR^2 + PR^3.... + PR^n\) we factorize out the common factor: \(A = PR(1+R + R^2 + R^3.... + PR^{n-1})\)

OpenStudy (anonymous):

Yes, correct. I have about four of these types of problems. So, if you give me the whole equation for this question (with the variables plugged in), I will be able to evaluate and solve the remaining three.

OpenStudy (anonymous):

596.10 + 596.10^2 + 596.10^3 .... etc?

OpenStudy (mathmate):

and use an identity that simplifies to: \(\huge A = PR\frac{R^n-1}{R-1}\) Note that this formula is different from the ones you have used/learned because she puts in he money at the BEGINNING of the period.

OpenStudy (mathmate):

Yes, you could do it the other way to check the answer as well! Good thought!

OpenStudy (anonymous):

SO... A = 596.10 * 21.10^14 - 1/21.10 - 1

OpenStudy (mathmate):

Your other way is not quite right, you need to do 585.551+596.294+607.238+....741.697 Make sure you calculate using at least 3 digits after the decimal to avoid rounding errors.

OpenStudy (mathmate):

Be careful with parentheses: \(\huge A = 575*1.01835 \frac{(1.01835^{14} - 1)}{(1.01835 - 1)}\)

OpenStudy (anonymous):

How did you get 1.01835?

OpenStudy (mathmate):

R=1+r/n=1+0.0365/2=1.01835 That is the equivalent cumulative rate of interest for every compounding period of six-months, including the principal.

OpenStudy (mathmate):

It's the same as the (1+r/n) in the compound interest formula, but makes the formula simpler.

OpenStudy (anonymous):

Ahhh.

OpenStudy (anonymous):

What about this one? The economy today grows at a rate of 2.2% semiannually. Ayden has an annuity that pays $1,655 at the beginning of every six month period. What is the value of Ayden’s annuity if he receives it in a lump sum now instead of over the next 18 years?

OpenStudy (mathmate):

Did you get the previous one, what did you get?

OpenStudy (anonymous):

9251.03

OpenStudy (mathmate):

Yes, that's what I got too! excellent.

OpenStudy (mathmate):

For this annuity problem, do you have the annuity formula? See if you can find it.

OpenStudy (anonymous):

PV of the annuity = 1655*(1 - 1.022^-36) / .022

OpenStudy (mathmate):

what would be the answer?

OpenStudy (anonymous):

40860.15833, but this is not an answer choice.

OpenStudy (anonymous):

P = PMT [(1 - (1 / (1 + r)n)) / r] Where: P = The present value of the annuity stream to be paid in the future PMT = The amount of each annuity payment r = The interest rate n = The number of periods over which payments are to be made

OpenStudy (mathmate):

That looks more like it. It's not the same as your previous.

OpenStudy (anonymous):

Ahh

OpenStudy (anonymous):

Hold on.

OpenStudy (mathmate):

wait, you used 1.022 instead of 1.011!

OpenStudy (mathmate):

I take it back, .022 is semi-annual interest.

OpenStudy (mathmate):

I'll do it myself and get back to you. Gimme a minute.

OpenStudy (mathmate):

It a bit similar to the previous problem because it's paid in advance (at the beginning of the period), so the future value is FV=1655*1.022(1.022^36-1)/0.022=91407.79 and convert it to PV PV=FV/(1.022^36)=41759.08 @mathater